Suppose we have 212 bytes of virtual memory and 27 bytes of physical main memory. Suppose the page size is 24 bytes.

a) How many pages are there in virtual memory?

b) How many page frames are there in main memory?

c) How many entries are in the page table for a process that uses all of virtual memory?
Answer #1

Virtual memory = 212 bytes

Physical memory = 27 bytes

Page size = 24 bytes

(a) No of pages in virtual memory = 212/24 = 8. I think virtual memory should be the exact multiple of page size.

(b) No of page frames = 27/24 = 1

(c) There is only one page frame in main memory. If a process uses all of the virtual memory, the process will be using 8 pages in virtual memory. So, there will be eight entries in its page table.